• A product has a price elasticity of supply of +3.0. Calculate how much quantity supplied changes if the price falls from $13 to $11.70, while output was originally 5,000 units.
1
Expert's answer
2021-12-09T15:47:48-0500

Solution:

Price elasticity of supply (PES) = "=\frac{\%\;change\; in\; quantity\; supplied}{\%\; change\; in\; price}" % change in quantity demanded/ % change in price

Let % change in quantity supplied = X

% change in price = "\frac{(11.70 - 13)}{(11.70 + 13)/2} = \frac{-1.3}{12.35} \times 100 = -10.53\%"


3.0 = "\frac{X}{-10.53\% }"

X = -10.53% "\times" 3.0 = -31.58"\%"

% Change in quantity demanded = -31.58%

Previous quantity = 5,000 units

New quantity = 5,000 – (5,000 "\times" -31.58%) = 3,421

Quantity supplied changed by = 5,000 – 3,421 = 1,579 units